A bag contains 12 bals of different colours: 5 red, 4 blue and 3 yellow. What is the probability of not selecting a red ball

The probability of an event not happening is 1 minus the probability of the event happening. Five of the twelve balls are red so the probability of selecting a red ball is 5/12. The probability of not selecting a red ball is 1 - 5/12. Find a common denominator in order to calculate this subtraction: 1 - 5/12 = 12/12 - 5/12 = 7/12.

LD

Related Maths GCSE answers

All answers ▸

factorise x^2 - 8x -65


A family go into a shop, they buy three sandwiches and two packets of crisps. It costs them £9. Another family buy five sandwiches and six packets of crisps. It costs them £19. How much does two sandwiches and five packets of crisps cost?


How can I find x and y?


solve the inequality x^2-4x+4>0